# Space-Time Graphs of Convex Sets for Multi-Robot Motion Planning

### v1

@@ I Introduction
Sampling-based planners, such as Rapidly-Exploring Random Trees (RRT) [1] and Probabilistic Roadmaps (PRM) [2], are popular for their simplicity and theoretical completeness. Yet spatiotemporal motion planning introduces additional challenges that significantly degrade their effectiveness: Dynamic obstacles can open or close narrow corridors, and random sampling may fail to capture these brief “windows” of safe transit. Moreover, many sampling-based planners either discretize time coarsely or incrementally adjust a time bound, limiting both effectiveness and efficiency in seeking time-optimal solutions with an unbounded time dimension. Finally, repeated collision checks, required for state expansions, become prohibitively expensive when the time dimension is included.
@@
Figure 1: Demonstration of the proposed PBS+TGCS method for MRMP: (a)(b) The solution visualized in (a) spatial coordinates and (b) space-time coordinates with the starts (squares), goals (stars), and trajectories (dot lines) of the robots highlighted in colors; (c) Low-level planner solving a TGCS program to generate a trajectory (blue solid line) through space-time collision-free convex sets (colored polyhedrons), which avoids trajectories of a high-priority robot (red solid line); (d) Magnified view of the congested region in (c), located at the center of the map.
Algorithmic Contributions: (1) We present the key idea of ST-GCS by demonstrating how to augment a space decomposition (i.e., a graph of convex sets) with an unbounded time dimension for time-optimal trajectories. By enforcing constraints on time flow and velocity bounds within a unified convex optimization, we obtain piecewise-linear, time-optimal trajectories without specialized time discretization. (2) We propose two MRMP methods by integrating ST-GCS into Random-Prioritized Planning (RP) [4], which randomly explores priority orders and plans each robot sequentially according to each priority order, and Priority-Based Search (PBS) [5], which searches over priority orders. In both frameworks, ST-GCS serves as the low-level trajectory planner once higher-priority trajectories are determined. (3) To make ST-GCS tractable when higher-priority trajectories are already planned, we introduce the Exact Convex Decomposition (ECD) algorithm to partition the space-time convex sets so that subsequent planning does not conflict with these “reserved” trajectories. Fig. 1(c)(d) illustrates how a robot’s trajectory is generated by applying ST-GCS on convex sets that have been updated to incorporate reserved trajectories as obstacle regions.
@@ II Related Work
Search-Based MAPF: MAPF [6] is a prominent approach for multi-robot path planning on graphs (e.g., 2D grids [7] or state lattices [8]) typically assuming discrete time steps. Modern search-based MAPF methods have offered powerful bi-level frameworks that decompose a multi-robot planning problem into the high-level coordination (e.g., PP [4], Conflict-Based Search [9], or PBS [5]) that resolves collisions among individual trajectories and low-level single-robot trajectory computations (e.g., space-time A[10]) that respects spatiotemporal constraints posed by the high level for collision resolution. Although some MAPF methods have been adapted to continuous-time robot actions [11, 12, 13], their solution quality remains constrained by the chosen discrete graph representation and limited motion primitives.
@@ III-A GCS: Motion Planning in Time-Invariant State Spaces
Following [29], we formulate the single-robot motion planning problem as a Mixed-Integer Convex Program (MICP) over $\mathcal{G}$. We introduce (1) a set of binary variables $\Phi=\{\phi_{e}\}_{e\in\mathcal{E}}$ that parametrizes any path $\pi_{\Phi}$, with $\phi_{e}=1$ iff $e\in\mathcal{E}(\pi_{\Phi})$; and (2) two vectors of continuous variables $\mathbf{x}_{v},\mathbf{y}_{v}\in\mathcal{X}_{v}$ that represent the robot’s initial and terminal states within each convex set $\mathcal{X}_{v},\forall v\in\mathcal{V}$. The MICP is:
@@
Here, $f(e)$ and $g(v)$ in Eqn. (1) specify additive costs over edges and vertices along the parametrized path $\pi_{\Phi}$. Given a feasible solution that fixes the values of $\pi_{\Phi}$, $\mathbf{x}$, and $\mathbf{y}$, we can reconstruct a continuous, collision-free trajectory $\tau$ by chaining the segments $(\mathbf{x}_{v},\mathbf{y}_{v})$ for each successive vertex in $\pi_{\Phi}$. Specifically, Constraints Eqn. (2) enforces that $\pi_{\Phi}$ is a simple path that uses edges in $\mathcal{(}E)$, which can be achieved by introducing auxiliary flow variables and constraints, commonly seen in linear program formulations for routing problems [31], rendering the above program nonlinear yet convex (see  [3] for more details). Constraints Eqn. (3) enforce that the two states $\mathbf{x}_{v},\mathbf{y}_{v}$ regarding each vertex $v\in\mathcal{V}$ must reside within the corresponding collision-free convex set $\mathcal{X}_{v}$, ensuring that the trajectory segment $(\mathbf{x}_{v},\mathbf{y}_{v})$ is also collision-free. Constraints Eqn. (4) enforce that the terminal state of $u$ always coincides with the initial state of $v$ for any edge $(u,v)$ of $\pi_{\Phi}$, ensuring that the reconstructed trajectory is continuous. Constraints Eqn. (5) ensure that $\tau$ starts from $\mathbf{x}_{\text{start}}$ and ends at $\mathbf{x}_{\text{goal}}$. Note that the two vertices (convex sets) $v_{\text{start}},v_{\text{goal}}\in\mathcal{V}$ are determined by iterating through all vertices $v\in\mathcal{V}$ to check whether $\mathbf{x}_{\text{start}}\in\mathcal{X}_{v}$ or $\mathbf{x}_{\text{goal}}\in\mathcal{X}_{v}$, respectively. [^1] In case of multiple $v_{\text{start}}$ (or $v_{\text{goal}}$), a hyper vertex is created connecting itself to each $v_{\text{start}}$ (or $v_{\text{goal}}$) with slight changes to constraints Eqn. (5).
@@ III-B ST-GCS: Time-Optimal Spatiotemporal Motion Planning
Although GCS effectively handles single-robot motion planning in a time-invariant $d$-dimansional state space, spatiotemporal motion planning with dynamic obstacles requires additional machinery to handle dynamic avoidance, variable arrival times, velocity bounds, and time optimality in a unified optimization. We thus propose ST-GCS, which operates on a graph of collision-free space-time convex sets.
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent a space-time decomposition, where each vertex $v\in\mathcal{V}$ corresponds to a space-time convex set $\mathcal{X}_{v}\subset\mathbb{R}^{d+1}$ that is free of both static and dynamic obstacles. We construct these sets by (1) extruding each given spatial convex set (free of static obstacles) from time 00 to a large finite $t_{\text{max}}$ [^2] In principle, time can remain unbounded. In practice, most MICP solvers require the convex sets to be finite and bounded. and (2) applying ECD (see Sec. IV-A) to remove space-time regions intersecting other robots’ trajectories (treated as dynamic obstacles), potentially subdividing these extruded sets further. An edge $(u,v)\in\mathcal{E}$ indicates $\mathcal{X}_{u}\cap\mathcal{X}_{v}\neq\emptyset$.
Figure 2: A continuous piecewise linear trajectory from $\mathbf{x}_{\text{start}}$ to $p_{\text{goal}}$ through collision-free 3D space-time convex sets (colored).
Let $\mathbf{x}.p$ and $\mathbf{x}.t$ respectively denote the spatial and temporal components of a space-time state $\mathbf{x}$. Within each space-time convex set $\mathcal{X}_{v}$, the local trajectory segment $\mathbf{x}_{v}\rightarrow\mathbf{y}_{v}$ is traversed at a uniform velocity $\mathbf{v}=\dfrac{\mathbf{y}_{v}.p-\mathbf{x}_{v}.p}{\mathbf{y}_{v}.t-\mathbf{x}_{v}.t}$, potentially different across sets. We now present the following ST-GCS formulation for spatiotemporal motion planning, which explicitly specifies a time-minimizing objective in Eqn. (6) and linear time and velocity constraints in Eqn. (8) and (9):
@@
s.t. Constraints in Eqn. (2)-(4) (7)
$\mathbf{y}_{v}.t-\mathbf{x}_{v}.t\geq\epsilon,$ $\forall v\in\pi_{\Phi}$ (8)
@@
Here, $\epsilon>0$ in Constraints Eqn. (8) is a small positive number to prevent time reversal, [^3] Eqn. (8) can be replaced by $\mathbf{y}_{v}.t-\mathbf{x}_{v}.t>0$, but most MICP solvers support only inclusive bounds. and Constraints Eqn. (9) impose given velocity bounds in each spatial dimension. Unlike Eqn. (5) in the static GCS formulation, Constraint Eqn. (10) only enforces that the spatial component matches the given goal position $p_{\text{goal}}$, leaving the arrival time unconstrained. Note that, while $v_{\text{start}}$ is determined in the same way as in GCS, each goal vertex $v\in\mathcal{V}_{\text{goal}}$ can be identified by checking whether $\mathcal{X}_{v}$ contains any states with the goal position (i.e., $\mathcal{X}_{v}\cap\{(p_{\text{goal}},t)\,|\,0\leq t\leq t_{\text{max}}\}\neq\emptyset$) and, if so, whether any arrival time at the goal position within $\mathcal{X}_{v}$ can be extended to $t_{\text{max}}$ (i.e., $\{(p_{\text{goal}},t)\,|\,t^{*}\leq t\leq t_{\text{max}}\}\subseteq\bigcup_{u\in\mathcal{V}}\mathcal{X}_{u}$, where $t^{*}=\min\{t\mid(p_{\text{goal}},t)\in\mathcal{X}_{v}\}$) to ensure the robot can stay there indefinitely.
As shown in Fig. 2, given a feasible solution that fixes the values of $\pi_{\Phi}$, $\mathbf{x}$, and $\mathbf{y}$, we can reconstruct a space-time piecewise linear trajectory, and any space-time state along the dashed line can serve as a valid goal if the robot can indefinitely stay at $p_{\text{goal}}$ the arrival. In summary, ST-GCS fuses the core convex path parametrization of GCS with explicit time and velocity constraints, enabling time-optimal, piecewise-linear trajectories in a space-time domain that can include dynamic obstacles (handled via ECD).
@@ IV ST-GCS For Multi-Robot Motion Planning
Figure 3: The proposed ECD algorithm decomposes (a) convex set $\mathcal{X}_{v_{i}}$ for a vertex $v_{i}\in\mathcal{V}$ and a trajectory segment $\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$ into (c)-(e) $6$ convex sets. (b) The trajectory segment is enlarged to a parallelepiped (highlighted in black) with its half-size being the safe radius between robots.
@@ IV-A Exact Convex Decomposition (ECD)
ECD takes as input a sequence of vertex-states tuples $\{(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{k}$, where each trajectory segment $\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$ lies entirely in a single convex set $\mathcal{X}_{v_{i}}$ of the given graph $\mathcal{G}$. We will discuss how to obtain such a sequence from any given piecewise linear trajectory resulting from an ST-GCS solution at the end of this subsection. Here, we present ECD for a 3D space-time state space, but the same principle can be extended to higher-dimensional Euclidian spaces with an additional time dimension.
Pseudocode (Alg. 1): ECD iterates over the $k$ vertex-states tuples [Line 1]. To simplify notation, let $\boxdot(p,r)$ denote the 2D square centered at $p$ with half-size $r$. For each tuple $(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})$, ECD first constructs a parallelepiped $L_{i}$ with its top and bottom faces being squares $\boxdot\left(\mathbf{x}_{i}.p,r\right)$ and $\boxdot\left(\mathbf{y}_{i}.p,r\right)$, respectively (see Fig. 3(b)). ECD then creates two convex sets $A_{i}$ and $B_{i}$ [Lines 1-1], bounding $L_{i}$ from above and below (see Fig. 3(b)). Specifically, $B_{0}$ excludes the cuboid formed by extruding $\boxdot\left(\mathbf{x}_{0}.p,r\right)$ from time 00 to $\mathbf{x}_{0}.t$, and $A_{k}$ excludes the cuboid formed by extruding $\boxdot\left(\mathbf{x}_{k}.p,r\right)$ from time $\mathbf{x}_{k}.t$ to $t_{\text{max}}$. This is crucial as it ensures the robot can safely stay at $\mathbf{x}_{0}.p$ during $t\in[0,\mathbf{x}_{0}.t)$ and at $\mathbf{x}_{k}.p$ during $t\in[\mathbf{x}_{k}.t,t_{\text{max}})$. As shown in Fig. 3(e), after removing $A_{i}\cup B_{i}$ from $\mathcal{X}_{v_{i}}$, ECD partitions the remaining center region into four convex sets by intersecting it with each halfspace $\text{outside}(L_{i},s)$ [Line 1], $s$ is each side face of $L_{i}$, considered as a separating plane, and $\text{outside}(L_{i},s)$ is the halfspace that does not contain $L_{i}$ (i.e., $\text{outside}(L_{i},s)\cap L_{i}=\emptyset$). ECD then removes $\text{outside}(L_{i},s)$ from the remaining region before processing the next halfspace [Line 1]. After partitioning $\mathcal{X}_{v_{i}}$ into these new convex sets, ECD adds a vertex corresponding to each new set to $\mathcal{G}$ [Line 1]. After processing all $k$ tuples, ECD updates the edges of $\mathcal{G}$ by checking intersections among new neighboring sets [Line 1]. Fig. 3 shows an example for an intermediate tuple $(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})$ with $1<i<k$.
Input: vertex-states tuples $\{(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{k}$, graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ of convex sets, safe radius $r$

### v2

@@ I Introduction
Sampling-based planners, such as Rapidly-Exploring Random Trees (RRT) [1] and Probabilistic Roadmaps (PRM) [2], are popular for their simplicity and theoretical completeness. However, spatiotemporal motion planning introduces additional challenges that significantly degrade their effectiveness: Dynamic obstacles can open or close narrow corridors, and random sampling may fail to capture these brief windows of safe transit. Moreover, many sampling-based planners either discretize time coarsely or incrementally adjust time bounds, limiting both effectiveness and efficiency in seeking time-optimal solutions within an arbitrarily large time dimension. Additionally, repeated collision checks required for state expansions become prohibitively expensive when the time dimension is included. When applied to MRMP in a coupled manner, sampling-based planners face even greater challenges due to the curse of dimensionality, which severely restricts scalability as the concatenated state space grows rapidly with the number of robots.
@@
Figure 1: Demonstration of the proposed PBS+ST-GCS for MRMP, where $R_{1}$ and $R_{3}$ exchange positions with $R_{2}$ and $R_{4}$, respectively. Left: Solution trajectories visualized in 2D spatial coordinates, with higher transparency indicating states at later time stamps. Right: Solution trajectories of $R_{1}$ and $R_{2}$ visualized in 3D space-time coordinates, where $R_{2}$ treats $R_{1}$ as a space-time obstacle and generates a trajectory through space-time collision-free convex sets (colored polyhedra) that exclude the trajectory of $R_{1}$.
Algorithmic Contributions: (1) We present the key idea of ST-GCS by demonstrating how to augment a spatial decomposition (i.e., a graph of convex sets) with an arbitrarily large time dimension for time-optimal trajectories. By enforcing constraints on time flow and velocity bounds within a unified convex optimization, we obtain piecewise-linear, time-optimal trajectories without specialized time discretization. (2) We propose two MRMP methods by integrating ST-GCS into Random-Prioritized Planning (RP) [4], which randomly explores priority orders and plans each robot sequentially according to each priority order, and Priority-Based Search (PBS) [5], which searches over priority orders. In both frameworks, ST-GCS serves as the low-level trajectory planner once higher-priority trajectories are determined. (3) To make ST-GCS tractable when higher-priority trajectories are already planned, we introduce the Exact Convex Decomposition (ECD) algorithm to partition the space-time convex sets so that subsequent planning does not conflict with these “reserved” trajectories. Fig. 1 illustrates how a robot’s trajectory is generated by applying ST-GCS on convex sets that have been updated to incorporate reserved trajectories as obstacle regions.
@@ II Related Work
Search-Based MAPF: MAPF [6, 7] is a prominent approach for multi-robot path planning on graphs (e.g., 2D grids [8] or state lattices [9]) typically assuming discrete time steps. Modern search-based MAPF methods have offered powerful bi-level frameworks that decompose a multi-robot planning problem into the high-level coordination (e.g., PP [4], Conflict-Based Search [10], or PBS [5]) that resolves collisions among individual trajectories and low-level single-robot trajectory computations (e.g., space-time A[11]) that respects spatiotemporal constraints posed by the high level for collision resolution. Although some MAPF methods have been adapted to continuous-time robot actions [12, 13, 14], their solution quality remains constrained by the chosen discrete graph representation and limited motion primitives.
@@ III-A GCS: Motion Planning in Time-Invariant State Spaces
Following [30], we formulate a bilinear program for the single-robot motion planning problem over $\mathcal{G}$, which is then transformed into a mixed-integer convex program for solving. We introduce (1) a set of binary variables $\Phi=\{\phi_{e}\}_{e\in\mathcal{E}}$ that parameterizes any path $\pi_{\Phi}$, with $\phi_{e}=1$ if and only if $e\in\mathcal{E}(\pi_{\Phi})$; and (2) two vectors of continuous variables $\mathbf{x}_{v},\mathbf{y}_{v}\in\mathcal{X}_{v}$ that represent the robot’s initial and terminal states within each convex set $\mathcal{X}_{v}$ for all $v\in\mathcal{V}$. For simplicity, we describe only the bilinear formulation below:
@@
where $f(e)$ and $g(v)$ in Eqn. (1) specify additive costs over edges and vertices along the parametrized path $\pi_{\Phi}$. Given a feasible solution that fixes the values of $\pi_{\Phi}$, $\mathbf{x}$, and $\mathbf{y}$, we can reconstruct a continuous, collision-free trajectory $\tau$ by chaining the segments $(\mathbf{x}_{v},\mathbf{y}_{v})$ for each successive vertex in $\pi_{\Phi}$. Specifically, Constraints Eqn. (2) enforces that $\pi_{\Phi}$ is a simple path that uses edges in $\mathcal{E}$, which can be achieved by introducing auxiliary flow conservation and degree constraints, commonly seen in linear program formulations for routing problems [32], rendering the above program nonlinear yet convex (see [3] for more details). Constraints Eqn. (3) enforce that the two states $\mathbf{x}_{v},\mathbf{y}_{v}$ regarding each vertex $v\in\mathcal{V}$ must reside within the corresponding collision-free convex set $\mathcal{X}_{v}$, ensuring that the trajectory segment $(\mathbf{x}_{v},\mathbf{y}_{v})$ is also collision-free. Constraints Eqn. (4) enforce that the terminal state of $u$ always coincides with the initial state of $v$ for any edge $(u,v)$ of $\pi_{\Phi}$, ensuring that the reconstructed trajectory is continuous. Constraints Eqn. (5) ensure that $\tau$ starts from $\mathbf{x}_{\text{start}}$ and ends at $\mathbf{x}_{\text{goal}}$. Note that the two vertices (convex sets) $v_{\text{start}},v_{\text{goal}}\in\mathcal{V}$ are determined by iterating through all vertices $v\in\mathcal{V}$ to check whether $\mathbf{x}_{\text{start}}\in\mathcal{X}_{v}$ or $\mathbf{x}_{\text{goal}}\in\mathcal{X}_{v}$, respectively. [^1] In case of multiple $v_{\text{start}}$ (or $v_{\text{goal}}$), a hyper vertex is created connecting itself to each $v_{\text{start}}$ (or $v_{\text{goal}}$[33] with slight changes to constraints Eqn. (5).
@@ III-B ST-GCS: Time-Optimal Spatiotemporal Motion Planning
Although GCS effectively handles single-robot motion planning in a time-invariant $d$-dimensional state space, spatiotemporal motion planning with dynamic obstacles requires additional machinery to handle dynamic avoidance, variable arrival times, velocity bounds, and time optimality in a unified optimization. We thus propose ST-GCS, which operates on a graph of collision-free space-time convex sets.
Figure 2: Piecewise linear trajectories from $\mathbf{x}_{\text{start}}$ to $p_{\text{goal}}$ through collision-free 2D space convex sets (lower-left) and 3D space-time convex sets (right).
Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ represent a space-time decomposition, where each vertex $v\in\mathcal{V}$ corresponds to a space-time convex set $\mathcal{X}_{v}\subset\mathbb{R}^{d+1}$ that is free of both static and dynamic obstacles. We construct these sets by (1) extruding each given spatial convex set (free of static obstacles) from time 00 to an arbitrarily large global time limit $t_{\text{max}}$ and (2) applying ECD (see Sec. IV-A) to remove space-time regions intersecting other robots’ trajectories (treated as dynamic obstacles), potentially subdividing the extruded sets further. An edge $(u,v)\in\mathcal{E}$ indicates $\mathcal{X}_{u}\cap\mathcal{X}_{v}\neq\emptyset$.
Let $\mathbf{x}.p$ and $\mathbf{x}.t$ respectively denote the spatial and temporal components of a space-time state $\mathbf{x}$. Within each space-time convex set $\mathcal{X}_{v}$, the local trajectory segment $\mathbf{x}_{v}\rightarrow\mathbf{y}_{v}$ is traversed at a uniform velocity $\mathbf{v}=\dfrac{\mathbf{y}_{v}.p-\mathbf{x}_{v}.p}{\mathbf{y}_{v}.t-\mathbf{x}_{v}.t}$, potentially different across sets. We now present the following ST-GCS formulation for spatiotemporal motion planning, which explicitly specifies a time-minimizing objective in Eqn. (6) and linear time and velocity constraints in Eqn. (8-9):
@@
s.t. Constraints in Eqn. (2-4) (7)
$\mathbf{y}_{v}.t-\mathbf{x}_{v}.t>0,$ $\forall v\in\pi_{\Phi}$ (8)
@@
where Constraints Eqn. (8) prevents time reversal from each state to a subsequent state  [^2] In practice, we impose Eqn. (8) as $\mathbf{y}_{v}.t-\mathbf{x}_{v}.t\geq\epsilon$ with an inclusive lower bound of a small positive number $\epsilon$.. Constraints Eqn. (9) impose given velocity bounds in each spatial dimension. Unlike Eqn. (5) in the static GCS formulation, Constraint Eqn. (10) only enforces that the spatial component matches the given goal position $p_{\text{goal}}$, leaving the arrival time unconstrained. Note that, while $v_{\text{start}}$ is determined in the same way as in GCS, each goal vertex $v\in\mathcal{V}_{\text{goal}}$ can be identified by checking whether $\mathcal{X}_{v}$ contains any states with the goal position (i.e., $\mathcal{X}_{v}\cap\{(p_{\text{goal}},t)\,|\,0\leq t\leq t_{\text{max}}\}\neq\emptyset$) and, if so, whether any arrival time at the goal position within $\mathcal{X}_{v}$ can be extended to $t_{\text{max}}$ (i.e., $\{(p_{\text{goal}},t)\,|\,t^{*}\leq t\leq t_{\text{max}}\}\subseteq\bigcup_{u\in\mathcal{V}}\mathcal{X}_{u}$, where $t^{*}=\min\{t\mid(p_{\text{goal}},t)\in\mathcal{X}_{v}\}$) to ensure the robot can stay there indefinitely.
As shown in Fig. 2, given a feasible solution that fixes the values of $\pi_{\Phi}$, $\mathbf{x}$, and $\mathbf{y}$, we can reconstruct a space-time piecewise linear trajectory, and any space-time state along the dashed line can serve as a valid goal if the robot can indefinitely stay at $p_{\text{goal}}$ the arrival. In summary, ST-GCS fuses the core convex path parametrization of GCS with explicit time and velocity constraints, enabling time-optimal  [^3] More rigorously, the time optimality is defined over the solution space of all piecewise linear trajectories parameterized by $(\Phi,\mathbf{x},\mathbf{y})$., piecewise-linear trajectories in a space-time domain that can include dynamic obstacles (handled via ECD).
@@ IV ST-GCS For Multi-Robot Motion Planning
Figure 3: The proposed ECD algorithm decomposes (a) convex set $\mathcal{X}_{v_{i}}$ for a vertex $v_{i}\in\mathcal{V}$ and a trajectory segment $\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$ into (c-e) $6$ convex sets. (b) The trajectory segment is enlarged to a parallelepiped (highlighted in black) with its apothem being the safe radius between robots.
@@ IV-A Exact Convex Decomposition (ECD)
ECD takes as input a sequence of vertex-segment tuples $\{(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{k}$, where each trajectory segment $\mathbf{x}_{i}\rightarrow\mathbf{y}_{i}$ lies entirely in a single convex set $\mathcal{X}_{v_{i}}$ of the given graph $\mathcal{G}$. We will discuss how to obtain such a sequence from any given piecewise linear trajectory resulting from an ST-GCS solution at the end of this subsection. Here, we assume each robot has a $d$-dimensional hypercube occupancy centered at each state $\mathbf{x}$ and exemplify ECD in a 3D space-time state space, but the same principle can be extended to other convex-shaped occupancies and to higher-dimensional state spaces with an additional time dimension.
Pseudocode (Alg. 1): ECD iterates over the $k$ vertex-segment tuples [Line 1]. To simplify notation, let $\boxdot(p,r)$ denote the 2D square centered at $p$ with apothem $r$. For each tuple $(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})$, ECD first constructs a parallelepiped $L_{i}$ with its top and bottom faces being squares $\boxdot\left(\mathbf{x}_{i}.p,r\right)$ and $\boxdot\left(\mathbf{y}_{i}.p,r\right)$, respectively (see Fig. 3(b)). ECD then creates two convex sets $A_{i}$ and $B_{i}$ [Lines 1-1], bounding $L_{i}$ from above and below (see Fig. 3(b)). Specifically, $B_{0}$ excludes the cuboid formed by extruding $\boxdot\left(\mathbf{x}_{0}.p,r\right)$ from time 00 to $\mathbf{x}_{0}.t$, and $A_{k}$ excludes the cuboid formed by extruding $\boxdot\left(\mathbf{x}_{k}.p,r\right)$ from time $\mathbf{x}_{k}.t$ to $t_{\text{max}}$. This is crucial as it ensures the robot can safely stay at $\mathbf{x}_{0}.p$ during $t\in[0,\mathbf{x}_{0}.t)$ and at $\mathbf{x}_{k}.p$ during $t\in[\mathbf{x}_{k}.t,t_{\text{max}})$. As shown in Fig. 3(e), after removing $A_{i}\cup B_{i}$ from $\mathcal{X}_{v_{i}}$, ECD partitions the remaining center region into four convex sets by intersecting it with each halfspace $\text{outside}(L_{i},s)$ [Line 1], where $s$ is each side face of $L_{i}$, considered as a separating plane, and $\text{outside}(L_{i},s)$ is the halfspace that does not contain $L_{i}$ (i.e., $\text{outside}(L_{i},s)\cap L_{i}=\emptyset$). ECD then removes $\text{outside}(L_{i},s)$ from the remaining region before processing the next halfspace [Line 1]. After partitioning $\mathcal{X}_{v_{i}}$ into these new convex sets, ECD adds a vertex corresponding to each new set to $\mathcal{G}$ [Line 1]. After processing all $k$ tuples, ECD updates the edges of $\mathcal{G}$ by checking intersections among new neighboring sets [Line 1]. Fig. 3 shows an example for an intermediate tuple $(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})$ with $1<i<k$.
Input: vertex-segment tuples $\{(v_{i},\mathbf{x}_{i},\mathbf{y}_{i})\}_{i=1}^{k}$, graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ of convex sets, safe radius $r$